A rectangle with an area of 300 square centimeters has a length and width in a ratio of 4:3. What are the length and width?
Answer by stanbon(75887)

A rectangle with an area of 300 square centimeters has a length and width in a ratio of 4:3. What are the length and width?
------
length = 4x
width = 3x
----
Equation:
4x*3x = 300 cm^2
----
12x^2 = 300
x^2 = 300/12
---
x = 5
------
length = 4x = 20 cm
width = 3x = 15 cm
-------------------------
Cheers,
Stan H.
---------------